# Investigating dusty Red Supergiant outflows in Westerlund 1 with 3D Hydrodynamic simulations

### v1

@@ 3.4 Dust Maps
We produce synthetic dust maps using the TORUS Monte Carlo radiative transfer code (Harries et al., 2019) with the procedure outlined in Mackey et al. (2021) to post-process our simulation. To account for the external radiation field that contributes to dust heating, we include the red supergiant itself as one source, and place two hot-star sources at positions cm and cm.
@@
The spectral energy distribution (SED) of the two hot-star sources are assumed to be blackbodies with an effective temperature kK, radius , and a luminosity . While this is more luminous than estimated values for individual stars in Westerlund 1 (the supergiants are at least a factor of three less luminous Clark et al., 2005; Beasor et al., 2021), the motivation for these parameters is to generate a radiation field comparable to that found in the inner regions of the Westerlund 1 cluster. In this way we avoid the complications of including tens of radiation sources with different positions and photospheric properties, making the simplest possible assumptions for this first study of the subject. At the location of the RSG, the resulting radiation energy flux from the two sources is 92.4 erg cm-2 s-1 and radiation energy density is eV cm-3. Further downstream the radiation flux decreases with the inverse square law, so the radiative heating is reduced.
@@
For the RSG we take a Kurucz (1992) stellar atmosphere model with K and stellar radius with a stellar mass of . This corresponds to a luminosity of . The RSG radiation field heats the dust in the immediate vicinity of the star, but in the downstream region the heating is dominated by the two hot radiation sources, particularly because the dust absorption cross-section peaks in the UV. We assume silicate grains (Draine, 2003) with a dust-to-gas ratio of 0.01, and a Mathis et al. (1977) grain-size distribution with power-law index -3.3 from 0.005 to 0.2 m. We further assume that grains do not exist in gas with temperatures K, i.e., instantaneous grain destruction. This is a gross simplification of the rich diversity of dust observed in winds of RSGs (e.g., Decin, 2021), but serves to obtain a reasonable estimate of the dust temperature and emissivity in the RSG wind. It would be relatively easy to obtain intensity variations of a factor of two or more by changing the dust composition, size distribution and dust-to-gas ratio within reasonable values, and so the predicted intensity maps have relatively large modelling uncertainties.
@@
We present synthetic dust continuum-emission maps for 11 m in Fig. 5 and 24 m in Fig. 6, as these are of interest to compare to current and possible future observations of RSG outflows in Westerlund 1 (Guarcello et al., 2024). The upper panel in each Fig.shows a projection with the line-of-sight being the -axis, and the lower panel shows the same but along the -axis. The logarithmic intensity scale brings out some of the faint features further in the downstream wake of the flow. The faintest features are starting to show some of the photon-sampling noise that is inherent in Monte-Carlo radiative transfer modelling.
@@
We see a range of clumps, globules with and without tails, and some filamentary emission connecting clumps that are being ablated from the RSG wind. These features are somewhat reminiscent of models of bow shocks from winds of cool runaway stars (Wareing et al., 2007; Mohamed et al., 2012), although here the much larger flow velocity of the cluster wind (and the higher resolution of our simulations) leads to a more highly clumped wake behind the star. The half opening-angle of the IR-emitting clumps in the downstream is about , with the majority of the dusty wind material enclosed within a cylinder of radius pc. It should be noted, however, that we impose a planar inflow rather than a spherically expanding cluster wind (which is relevant on scales of a few parsecs) and so we probably underestimate the dispersion of the clumps in directions perpendicular to the cluster-wind flow ().
Projections of the 11 m and 24 m emission from different viewing perspectives are shown in Fig. 7, progressing from face-on (, where is the angle between the line-of-sight and the cluster-wind flow velocity) to edge-on () from left to right, respectively. The face-on images have a circular symmetry with a ‘fried-egg’ morphology that results from a peak of emission in the termination-shock of the RSG wind. As increases, the nebula acquires the head-tail morphology of an ablation flow, with dynamical instabilities and a turbulent wake. At 11 m, Fig. 7 shows an emission feature that resembles a bow shock from the edge-on perspective, although the the cluster wind is only transsonic and no emission from it is visible in dust maps. This feature is the shocked RSG wind, which accumulates in a dense isothermal layer behind the RSG wind termination-shock and then advects downstream around the unshocked RSG wind region.
@@ 3.5 Comparison to observations
We compare our synthetic observations with the JWST/MIRI F1130W observations towards Westerlund 1 by Guarcello et al. (2024), taken as part of GO-1905 (PI Guarcello). Since our objective is a phenomenological comparison rather than bespoke modelling of a specific system, we use the level 3 data products from the MAST111https://archive.stsci.edu/ science archive without further processing or reduction.
@@
We focus on three M-type supergiants: W20, W26 and W237, which are plotted in Fig.8. In each case there is a substantial saturated inner zone, simply due to the high sensitivity of JWST and bright nature of the sources. However, exterior to that is a plethora of unsaturated features.
@@
Qualitatively and quantitatively, the 11 m synthetic image from Fig. 7 at is very similar to the JWST F1130W image of the RSG W237 (Guarcello et al., 2024, see also Fig. 8). The synthetic images at and are also remarkably similar to the F1130W image of the RSG W20, although there the JWST image seems to show brighter emission that our prediction. For both W20 and W26, a detailed comparison with simulations is difficult because large parts of the nebula are saturated in the F1130W image.
@@
The synthetic 24 m emission maps in Fig. 7 show emission that is more than brighter than at 11 m, as discussed already for Fig. 6. Plotting with a linear rather than logarithmic intensity scale brings out the chaotic and filamentary morphology of the wake behind the star, with clumps connected by ridges of dusty gas as the ablation process rips them away from the bulk of the RSG wind material.
@@
The fluxes in the unsaturated parts of the extended features associated with the M-type supergiants are typically in the range 3000-8000 MJy ster-1, comparable to (i.e., same order of magnitude as) the values in the simulated maps of Figures 5, 6, and 7. The morphology of the flow is also similar, with a primary flow, multiple cometary secondary objects and even extended features that connect to the primary flow that likely result from some combination of Rayleigh-Taylor/Kelvin-Helmholtz instabilities, as illustrated in Fig.9. Such features have previously been discussed in the context of runaway AGB stars such as Mira (Wareing et al., 2007), although here the numerical resolution is significantly higher and a more complex flow structure is apparent.
@@
It is notable that the model predicts similar fluxes to the observations without any contribution from polycyclic aromatic hydrocarbon (PAH) emission, despite there being a significant PAH C-H bending feature within this band (e.g., Tappe et al., 2006; Candian & Sarre, 2015; Wesson et al., 2024). This suggests that the PAH abundance in the ablation flow is depleted, most likely due to destruction by shocks or radiation (e.g., Micelotta et al., 2010; Siebenmorgen & Krügel, 2010). This appears to be supported by ALMA continuum observations of free-free emission from the nebulae around the cool supergiants in Westerlund 1 (Fenech et al., 2018), ATCA radio observations that also detect extended nebulae around the cool supergiants (Andrews et al., 2018), and H emission from the nebula around W26 (Wright et al., 2014), all implying that the winds of the cool supergiants are externally photoionized. However, confirming a deficit of PAH abundance will require a multi-wavelength study with dedicated models that treat PAHs, which is beyond the scope of this work.
@@ 4 Discussion
4.1 RSG & Cluster Wind Uncertainties
The mass-loss rates of RSGs are poorly constrained (Beasor et al., 2020) and can vary rapidly over time, including strongly clumped mass-loss episodes (Montargès et al., 2021). In this work, we chose a constant “canonical” value, but a time-varying mass-loss rate would increase the stochasticity of the mass-loading rate, as well as the morphology of dust emission. The cluster wind properties are also poorly constrained from observations. To fully account for this, a model of the cluster wind’s velocity, density, and temperature as a function of radius would be required. We also assume a planar inflow here, whereas in reality it is spherically expanding, with potentially strong local perturbations from nearby individual stars. The cluster wind expansion radius is on the order of the cluster radius itself, and therefore could impact the cluster wind at the positions of the RSGs. We see the effects of our initial conditions for the first 25-50 kyr of our simulation, as we “switch on” our star. In reality, the RSG would have evolved from a main sequence star with a hot, fast wind. The RSG would expand into the bubble excavated during the main sequence as opposed to the uniform ISM we assume here.
@@ 4.2 Modelling Uncertainties
Many previous works have focused on the physics of radiative turbulent mixing layers, which arise in many different astrophysical contexts (e.g., Fielding et al., 2020; Tan et al., 2021; Mackey et al., 2025). Fielding et al. (2020) showed that while the phase structure is resolution dependent, the total cooling is well converged even for moderate resolutions. We do not consider the effects of magnetic fields in this work. We would expect the morphology of the dust filaments to be elongated along the cluster magnetic field lines (Gronke & Oh, 2020) but the contribution to the turbulent mixing is likely not significant (Li et al., 2020). We also do not consider thermal conduction in this work. The importance of thermal conduction is debated in the literature, but recent work suggests that its effects are only significant if larger than the turbulent diffusivity (Ji et al., 2019; Tan et al., 2021).
Our dust modelling has several uncertainties. We assume a generic hot radiation field, but the dust emission is sensitive to the specifics of this radiation field. For example, a hot star in the negative- direction would heat up the dust and change the observed dust map considerably. We also make assumptions about the properties of the dust grains, and the ratio of gas to dust in the ablation flow. Given the number of uncertainties and the fact that we did not tune any parameters, the qualitative and quantitative agreement between our synthetic 11m emission maps and the observed emission around the RSGs in Westerlund 1 is encouraging.
@@ 4.3 Mass loading
We obtain a lower limit for mass-loading efficiency due to RSG wind material being accelerated and exiting the domain before it can be heated, as well as due to resolution as we discuss in Appendix A. Using the cluster wind velocity formalism given in Eq. 1 of Härer et al. (2023), and assuming a total current RSG mass-loss rate in Westerlund 1 of , we would expect the cluster wind to be slowed down by a factor due to RSG winds. Westerlund 1 is a young cluster with a small number of RSGs at present, so here this contribution is probably not important. At later times however, when a cluster wind is weaker due to fewer WR stars and mass loading is stronger due to more RSG stars, this effect could become more important. Furthermore, the JWST observations (Guarcello et al., 2024) show that there is significant mid-IR nebular emission in all directions around the star cluster, some of which cannot have come from the RSGs unless the cluster-wind geometry is very unusual. This material also contributes to mass-loading of the cluster wind and so the combined effects of all of the cool, dusty gas is larger than the 10% effect we estimate for just the RSG winds.
@@
In this work we only consider a single set of steady-state stellar wind parameters, and future work is required to examine the effects of initial stellar mass and stellar evolution on mass-loading in clusters. We also neglect other possibilities such as bursts and asymmetries in mass-loss, which would affect the morphology of the ablation structures and dust emission.
4.4 Relevance to SSCs
@@
Super star-clusters are extremely massive YMCs that are sufficiently massive and tightly bound that they could be progenitors of globular clusters, which are often observed to have multiple stellar populations (Gratton et al., 2012; Bastian & Lardo, 2018). The enrichment patterns of the second-generation stars imply that they formed from gas that was polluted by nuclear-processed gas ejected from the first generation of (massive) stars. Feedback from (hot) massive stars is one explanation for these multiple populations (e.g., Wünsch et al., 2008; de Mink et al., 2009; Krause et al., 2013), and cool supergiants have also been proposed (e.g. Szécsi et al., 2018). Rather specific conditions are required in order for winds of hot stars to cool efficiently and remain bound to the star cluster to form a second generation (Wünsch et al., 2017). RSG winds are already cool, dense and dusty, and thus readily suited to forming a second generation of stars if the wind material can be confined within the cluster. While Westerlund 1 may not be massive enough to retain its gas and form a 2nd generation in situ, it is very interesting to study (presumably) cool and dusty gas right in the core of the cluster. Measuring the thermal and chemical properties of this gas, and estimating its lifetime in the cluster, may give insights into how the multiple populations in globular clusters could have formed. This is especially interesting because some targeted searches for intra-cluster gas and dust failed to detect either in significant quantities in other star clusters (Bastian & Strader, 2014).
4.5 Follow up observations
@@
Given the high velocities associated with the ablation flow, of order 1000 km s-1, they should be spectrally resolvable even at medium resolution. VLT/ERIS-SPIFFER (Davies et al., 2023) is a near-infrared integral field spectrograph facility with a high resolution K-middle filter including Br (e.g. Reiter et al., 2024). This should therefore be well suited to confirming those high velocities without the same susceptibility to saturation that JWST suffers. That would immediately rule out a primordial origin of the nebular emission seen towards Westerlund 1 and confirm that it instead results from stellar ejecta. It would also provide the ability to make empirical mass-loss rate estimates, which is a key factor in interpreting these ablation flows.
@@
Westmoquette et al. (2010) and Westmoquette et al. (2013) obtained spectra of dynamical mixing layers at the boundary between hot cluster winds and pillars of dense gas and dust. They could measure the enhanced velocity dispersion in the mixing layer of about 100 km s-1. Similar observations of these dusty clumps in Westerlund 1 should give observational constraints on the degree of mixing between hot and cool phases, and hence on the mass-loading of the cluster wind. Comparison of such observations with synthetic spectra from hydrodynamical simulations will be valuable for constraining physical processes operating in the mixing layer (e.g., hydrodynamical mixing vs. thermal conduction).

### v2

@@ 3.4 Dust maps
We produced synthetic dust maps using the TORUS Monte Carlo radiative transfer code (Harries et al., 2019) with the procedure outlined in Mackey et al. (2021) to post-process our simulation. To account for the external radiation field that contributes to dust heating, we included the RSG itself as one source, and placed two hot-star sources at positions cm and cm.
@@
The spectral energy distribution (SED) of the two hot-star sources are assumed to be blackbodies with an effective temperature kK, radius , and a luminosity . While this is more luminous than estimated values for individual stars in Westerlund 1 (the supergiants are at least a factor of three less luminous; Clark et al., 2005; Beasor et al., 2021), the motivation for these parameters is to generate a radiation field comparable to that found in the inner regions of the Westerlund 1 cluster. In this way we avoid the complications of including tens of radiation sources with different positions and photospheric properties, making the simplest possible assumptions for this first study of the subject. At the location of the RSG, the resulting radiation energy flux from the two sources is 92.4 erg cm-2 s-1 and radiation energy density is eV cm-3. Further downstream the radiation flux decreases with the inverse square law, so the radiative heating is reduced.
@@
For the RSG we took a Kurucz (1992) stellar atmosphere model with K and stellar radius with a stellar mass of . This corresponds to a luminosity of . The RSG radiation field heats the dust in the immediate vicinity of the star, but in the downstream region the heating is dominated by the two hot radiation sources, particularly because the dust absorption cross-section peaks in the UV. We assumed silicate grains (Draine, 2003) with a dust-to-gas ratio of 0.01, and a Mathis et al. (1977) grain-size distribution with power-law index of -3.3 from 0.005 to 0.2 m. We further assumed that grains do not exist in gas with temperatures K, i.e. instantaneous grain destruction. This is a gross simplification of the rich diversity of dust observed in winds of RSGs (e.g. Decin, 2021), but serves to obtain a reasonable estimate of the dust temperature and emissivity in the RSG wind. It would be relatively easy to obtain intensity variations of a factor of two or more by changing the dust composition, size distribution and dust-to-gas ratio within reasonable values, and so the predicted intensity maps have relatively large modelling uncertainties.
@@
We present synthetic dust continuum-emission maps for 11 m in Fig. 5 and 24 m in Fig. 6, as these are of interest to compare to current and possible future observations of RSG outflows in Westerlund 1 (Guarcello et al., 2025). The upper panel in each figure shows a projection with the line of sight being the -axis, and the lower panel shows the same but along the -axis. The logarithmic intensity scale brings out some of the faint features further in the downstream wake of the flow. The faintest features are starting to show some of the photon-sampling noise that is inherent in Monte Carlo radiative transfer modelling.
@@
We see a range of clumps, globules with and without tails, and some filamentary emission connecting clumps that are being ablated from the RSG wind. These features are somewhat reminiscent of models of bow shocks from winds of cool runaway stars (Wareing et al., 2007; Mohamed et al., 2012), although here the much larger flow velocity of the cluster wind (and the higher resolution of our simulations) leads to a more highly clumped wake behind the star. The half opening-angle of the IR-emitting clumps in the downstream is about , with the majority of the dusty wind material enclosed within a cylinder of radius pc. It should be noted, however, that we imposed a planar inflow rather than a spherically expanding cluster wind (which is relevant on scales of a few parsecs), and so we probably underestimate the dispersion of the clumps in directions perpendicular to the cluster-wind flow (). The optical depth of the clumps to EUV/far-UV and optical stellar radiation is (depending on clump size and photon energy), calculated using the opacity tables from torus, and so they are not self-shielding and are photoionised throughout to a high degree.
Projections of the 11 m and 24 m emission from different viewing perspectives are shown in Fig. 7, progressing from face-on (, where is the angle between the line of sight and the cluster-wind flow velocity) to edge-on () from left to right. The face-on images have a circular symmetry with a ‘fried-egg’ morphology that results from a peak of emission in the termination-shock of the RSG wind. As increases, the nebula acquires the head-tail morphology of an ablation flow, with dynamical instabilities and a turbulent wake. At 11 m, Fig. 7 shows an emission feature that resembles a bow shock from the edge-on perspective, although the cluster wind is only trans-sonic and no emission from it is visible in dust maps. This feature is the shocked RSG wind, which accumulates in a dense isothermal layer behind the RSG wind termination-shock and then advects downstream around the unshocked RSG wind region.
@@ 3.5 Comparison to observations
We compared our synthetic observations with the JWST/MIRI F1130W observations towards Westerlund 1 by Guarcello et al. (2025), taken as part of GO-1905 (PI Guarcello). Since our objective is a phenomenological comparison rather than bespoke modelling of a specific system, we used the level 3 data products from the MAST111https://archive.stsci.edu/ science archive without further processing or reduction.
@@
We focused on three M-type supergiants: W20, W26, and W237, which are plotted in Fig.8. In each case there is a substantial saturated inner zone, simply due to the high sensitivity of JWST and bright nature of the sources. However, exterior to that is a plethora of unsaturated features.
@@
Qualitatively and quantitatively, the 11 m synthetic image from Fig. 7 at is very similar to the JWST F1130W image of the RSG W237 (Guarcello et al. 2025; see also our Fig. 8). The synthetic images at and are also remarkably similar to the F1130W image of the RSG W20, although there the JWST image seems to show brighter emission that our prediction. For both W20 and W26, a detailed comparison with simulations is difficult because large parts of the nebula are saturated in the F1130W image.
@@
The synthetic 24 m emission maps in Fig. 7 show emission that is more than ten times brighter than at 11 m, as discussed already for Fig. 6. Plotting with a linear rather than logarithmic intensity scale brings out the chaotic and filamentary morphology of the wake behind the star, with clumps connected by ridges of dusty gas as the ablation process rips them away from the bulk of the RSG wind material.
@@
The fluxes in the unsaturated parts of the extended features associated with the M-type supergiants are typically in the range 3000-8000 MJy ster-1, comparable to (i.e. same order of magnitude as) the values in the simulated maps of Figs. 5, 6, and 7. The morphology of the flow is also similar, with a primary flow, multiple cometary secondary objects and even extended features that connect to the primary flow that likely result from some combination of RT and/or KH instabilities, as illustrated in Fig.9. Such features have previously been discussed in the context of runaway asymptotic giant branch stars such as Mira (Wareing et al., 2007), although here the numerical resolution is significantly higher and a more complex flow structure is apparent.
@@
It is notable that the model predicts similar fluxes to the observations without any contribution from polycyclic aromatic hydrocarbon (PAH) emission, despite there being a significant PAH C-H bending feature within this band (e.g. Tappe et al., 2006; Candian & Sarre, 2015; Wesson et al., 2024). This suggests that the PAH abundance in the ablation flow is depleted, most likely due to destruction by shocks or radiation (e.g. Micelotta et al., 2010; Siebenmorgen & Krügel, 2010). We note that PAHs are not universally present in RSG winds (Verhoelst et al., 2009) and so it is possible that they are not formed in the first place. Future observations may be able to resolve this question.
This appears to be supported by ALMA continuum observations of free-free emission from the nebulae around the cool supergiants in Westerlund 1 (Fenech et al., 2018), ATCA radio observations that also detect extended nebulae around the cool supergiants (Andrews et al., 2018), and H emission from the nebula around W26 (Wright et al., 2014), all implying that the winds of the cool supergiants are externally photoionised. However, confirming a deficit of PAH abundance will require a multi-wavelength study with dedicated models that treat PAHs, which is beyond the scope of this work.
@@ 4 Discussion
4.1 RSG and cluster wind uncertainties
The mass-loss rates of RSGs are poorly constrained (Beasor et al., 2020) and can vary rapidly over time, including strongly clumped mass-loss episodes (Montargès et al., 2021). In this work we chose a constant ‘canonical’ value, but a time-varying mass-loss rate would increase the variation of the mass-loading rate and shedding timescale, as well as the morphology of dust emission. The cluster wind properties are also poorly constrained from observations. To fully account for this, a model of the cluster wind’s velocity, density, and temperature as a function of radius would be required. We also assumed a planar inflow here, whereas in reality it is spherically expanding, with potentially strong local perturbations from nearby individual stars. The cluster wind expansion radius is of the order of the cluster radius itself, and therefore could impact the cluster wind at the positions of the RSGs. We see the effects of our initial conditions for the first 25-50 kyr of our simulation, as we ‘switch on’ our star. In reality, the RSG would have evolved from a main sequence star with a hot, fast wind. The RSG would expand into the bubble excavated during the main sequence as opposed to the uniform ISM we assumed here.
@@ 4.2 Modelling uncertainties
Many previous works have focused on the physics of radiative turbulent mixing layers, which arise in many different astrophysical contexts (e.g. Fielding et al., 2020; Tan et al., 2021; Mackey et al., 2025). Fielding et al. (2020) show that while the phase structure is resolution dependent, the total cooling is well converged even for moderate resolutions. We did not consider the effects of magnetic fields in this work. We would expect the morphology of the dust filaments to be elongated along the cluster magnetic field lines (Gronke & Oh, 2020) but the contribution to the turbulent mixing is likely not significant (Li et al., 2020). We did not consider thermal conduction in this work either. The importance of thermal conduction is debated in the literature, but recent work suggests that its effects are only significant if larger than the turbulent diffusivity (Ji et al., 2019; Tan et al., 2021).
Our dust modelling has several uncertainties. We assumed a generic hot radiation field, but the dust emission is sensitive to the specifics of this radiation field. For example, a hot star in the negative- direction would heat up the dust and change the observed dust map considerably. We also made assumptions about the properties of the dust grains, and the ratio of gas to dust in the ablation flow. Given the number of uncertainties and the fact that we did not tune any parameters, the qualitative and quantitative agreement between our synthetic 11m emission maps and the observed emission around the RSGs in Westerlund 1 is encouraging.
@@ 4.3 Mass loading
We obtain a lower limit for mass-loading efficiency due to RSG wind material being accelerated and exiting the domain before it can be heated, as well as due to resolution as we discuss in Appendix A. Using the cluster wind velocity formalism given in Eq. 1 of Härer et al. (2023), and assuming a total current RSG mass-loss rate in Westerlund 1 of -1 , we would expect the cluster wind to be slowed by a factor of due to RSG winds. Westerlund 1 is a young cluster with a small number of RSGs at present, so here this contribution is probably not important. At later times however, when a cluster wind is weaker due to fewer WR stars and mass loading is stronger due to more RSG stars, this effect could become more important. Furthermore, the JWST observations (Guarcello et al., 2025) show that there is significant mid-IR nebular emission in all directions around the star cluster, some of which cannot have come from the RSGs unless the cluster-wind geometry is very unusual. This material also contributes to the mass loading of the cluster wind, and so the combined effects of all of the cool, dusty gas is larger than the 10% effect we estimate for just the RSG winds.
@@
In this work we only considered a single set of steady-state stellar wind parameters, and future work is required to examine the effects of initial stellar mass and stellar evolution on mass loading in clusters. We also neglected other possibilities such as bursts and asymmetries in mass-loss, which would affect the morphology of the ablation structures and dust emission.
4.4 Relevance to super star clusters
@@
Super star-clusters are extremely massive YMCs that are sufficiently massive and tightly bound that they could be progenitors of globular clusters, which are often observed to have multiple stellar populations (Gratton et al., 2012; Bastian & Lardo, 2018). The enrichment patterns of the second-generation stars imply that they formed from gas that was polluted by nuclear-processed gas ejected from the first generation of (massive) stars. Feedback from (hot) massive stars is one explanation for these multiple populations (e.g. Wünsch et al., 2008; de Mink et al., 2009; Krause et al., 2013), and cool supergiants have also been proposed (e.g. Szécsi et al., 2018). Rather specific conditions are required in order for winds of hot stars to cool efficiently and remain bound to the star cluster to form a second generation (Wünsch et al., 2017). RSG winds are already cool, dense and dusty, and thus readily suited to forming a second generation of stars if the wind material can be confined within the cluster. While Westerlund 1 may not be massive enough to retain its gas and form a 2nd generation in situ, it is very interesting to study (presumably) cool and dusty gas right in the core of the cluster. Measuring the thermal and chemical properties of this gas, and estimating its lifetime in the cluster, may give insights into how the multiple populations in globular clusters could have formed. This is especially interesting because some targeted searches for intra-cluster gas and dust failed to detect either in significant quantities in other star clusters (Bastian & Strader, 2014).
4.5 Follow-up observations
@@
Given the high velocities associated with the ablation flow, of order 1000 km s-1, they should be spectrally resolvable even at medium resolution. VLT/ERIS-SPIFFER (Davies et al., 2023) is a near-IR integral field spectrograph facility with a high resolution K-middle filter including Br (e.g. Reiter et al., 2024). This should therefore be well suited to confirming those high velocities without the same susceptibility to saturation that JWST suffers. That would immediately rule out a primordial origin for the nebular emission seen towards Westerlund 1 and confirm that it instead results from stellar ejecta. It would also provide the ability to make empirical mass-loss rate estimates, which is a key factor in interpreting these ablation flows.
@@
Westmoquette et al. (2010, 2013) obtained spectra of dynamical mixing layers at the boundary between hot cluster winds and pillars of dense gas and dust. They were able to measure the enhanced velocity dispersion in the mixing layer of about 100 km s-1. Similar observations of these dusty clumps in Westerlund 1 should give observational constraints on the degree of mixing between hot and cool phases, and hence on the mass loading of the cluster wind. Comparison of such observations with synthetic spectra from HD simulations will be valuable for constraining physical processes operating in the mixing layer (e.g. HD mixing vs thermal conduction).